A man is running with 6m/s with his umbrella held vertically. If he doubles his speed he holds the umbrella at 30 degree with vertical.find the speed and direction of rain


1
Expert's answer
2021-07-12T11:32:16-0400



"v_m=2\times 6=12ms^{-1}\\
\tan30\degree=\dfrac{12}{v_{rm}}\\
\\
{v_{rm}}=12\sqrt{3}ms^{-1}\\
v_r=\sqrt{v_m^2+v_{rm}^2}\\
v_r=\sqrt{12^2+(12\sqrt{3})^2}=24ms^{-1}\\
\theta+30\degree=90\degree\\
\theta=60\degree\\"

Rain speed is 24 metres per second in the direction 60 degrees East of South